OpenStudy (anonymous):

Will medal: Factor the trinomial x^2-18x+81 a: (x-8)^2 b: (x-9)^2 c: (x+8)^2 d: (x+9)^2 e: none of these

OpenStudy (math&ing001):

x^2 - 18x + 81 = x^2 - 2*9*x + 9^2

OpenStudy (math&ing001):

a^2 - 2*b*a + b^2 = (a-b)^2